OpenStudy (anonymous):

Choose one of the factors of x3 – 1331 x – 11 x2 – 11x + 121 x2 + 22x + 121 None of the above This is about the sum or difference in cubs (factoring) and I'm pretty sure it's "none of the above." could anyone double check?

OpenStudy (anonymous):

cubes*

OpenStudy (lgbasallote):

hint 1331 = 11^3

OpenStudy (anonymous):

I did the work already, but think it's wrong @lgba.

OpenStudy (lgbasallote):

therefore you can use difference of two cubes a^3 - b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (lgbasallote):

a = x and b = 11

OpenStudy (anonymous):

(x-11)(x^2+11x+121) Would this be correct?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

so is the answer obvious now? :D

OpenStudy (anonymous):

Oh okay tyvm, I didn't know it the addition signs in the 2nd one would turn to subtraction cause of the -11. Thx :) and yeah it's really easy :) All you need to know is the formula thingie you posted above :)

OpenStudy (lgbasallote):

yup..the formula for cubes is sometimes tricky in my opinion haha

OpenStudy (anonymous):

i still dont get it